### v1

# TikTokActions: A TikTok-Derived Video Dataset for Human Action Recognition

Abstract
The increasing variety and quantity of tagged multimedia content on platforms such as TikTok provides an opportunity to advance computer vision modeling. We have curated a distinctive dataset of 283,582 unique video clips categorized under 386 hashtags relating to modern human actions. We release this dataset as a valuable resource for building domain-specific foundation models for human movement modeling tasks such as action recognition. To validate this dataset, which we name TikTokActions, we perform two sets of experiments. First, we pretrain the state-of-the-art VideoMAEv2 with a ViT-base backbone on TikTokActions subset, and then fine-tune and evaluate on popular datasets such as UCF101 and the HMDB51. We find that the performance of the model pre-trained using our TikTok dataset is comparable to models trained on larger action recognition datasets(95.3% on UCF101 and 53.24% on HMDB51). Furthermore, our investigation into the relationship between pre-training dataset size and fine-tuning performance reveals that beyond a certain threshold, the incremental benefit of larger training sets diminishes. This work introduces a useful TikTok video dataset that is available for public use and provides insights into the marginal benefit of increasing pre-training dataset sizes for video-based foundation models.
1 Introduction
Action recognition is a fundamental computer vision task that involves identifying and classifying human actions in video sequences, with significant applications in areas such as surveillance, interactive gaming, and healthcare. In the era of social media platforms, TikTok has emerged as a significant repository of human action videos. We have compiled a dataset of 283,582 TikTok video clips that encompass a range of human activities, including dance, sports, fitness, hand gestures, and daily routines. This collection aims to represent various human actions in a manner that is relevant to real-world scenarios and cultural contexts. Importantly, this dataset is uniquely curated to include well-defined but non-standard human actions, filling gaps in existing action recognition datasets and diversifying the range of human activities recognized by current models. We developed a model for human action recognition trained using the curated TikTok video clips. Pre-training on a dataset with varied and authentic content is intended to more accurately reflect the spectrum of naturalistic human actions, aiming to address the limitations of conventional datasets by incorporating elements of contemporary behavior. In our initial evaluations, presented later in this paper, we find that utilizing our TikTok dataset may enhance the development of targeted human action recognition models in the field of computer vision.
Moreover, the nature of our data source, TikTok, presents another advantage. Kinetics[4] derives its content from YouTube, featuring a combination of professional and amateur videos, whereas TikTok’s database consists primarily of user-generated material. This means that our dataset is inherently rich with real-world scenarios: from spontaneous camera movements and varied illumination conditions to diverse backgrounds and authentic human interactions. This rawness provides a genuine challenge and opportunity for training robust models capable of understanding human actions in a plethora of real-world conditions.
We conducted a series of experiments to assess the quality of our dataset and its potential to serve as a basis for pre-training foundation models for action-related tasks. We used a well-established state-of-the-art model VideoMAE V2[9], and pre-trained it on our TikTokActions dataset. We then fine-tuned the model on two well-known video action recognition datasets: UCF101[7] and HMDB51[5]. Our experiments provided valuable insights: the performance achieved was comparable to models that were both pre-trained and fine-tuned solely on UCF101 and HMDB51. This underscores the utility of the TikTokActions dataset to serve as a useful pre-training dataset for action-related tasks.
We next delved into understanding the relationship between the quantity of pre-training videos and the performance on the downstream action recognition task. By sampling subsets of our TikTok dataset, starting with 1,000 videos and increasing in increments of 1,000 up to 6,000 videos, pre-training VideoMAE V2[9] on the sampled subset, and subsequently fine-tuning each on the benchmark datasets, we observed that as long as the number of pre-training videos remains above a certain threshold, the performance remains comparable to massive pre-training datasets. This provides a valuable lesson when pre-training foundation models: the benefits of increasing the pre-training dataset size diminish as the dataset increases.
2 Related Work
2.1 Human Action Recognition Datasets
The evolution of action recognition research has been strongly influenced by the availability of benchmark datasets. The UCF101[7] dataset, with its diverse collection of human actions, has been a cornerstone for early research in this domain, guiding and benchmarking numerous computer vision innovations applied to action recognition. Similarly, the HMDB-51[5] dataset, despite its limited size, has been pivotal in providing a varied set of human actions captured under realistic conditions.
The Kinetics Human Action Video dataset[4], introduced in 2017, marked a significant milestone in this field. With its vast collection of YouTube-sourced videos encompassing 400 distinct human action classes, Kinetics has become an indispensable resource for researchers, setting a higher benchmark and promoting advanced research in action recognition.
Another noteworthy dataset is the Something-Something[3] collection. Its unique focus on human-object interactions captured in real-world scenarios has offered researchers a different perspective, emphasizing the complexity and diversity of everyday actions.
These datasets, along with several others, have collectively laid the foundation for the current advancements in action recognition, facilitating the development and evaluation of state-of-the-art models and techniques. Recognizing the rapid evolution of content on platforms like TikTok and the emergence of diverse, contemporary actions and scenarios, we are motivated to develop a comprehensive dataset highlighting the latest trends and user behaviors specific to new short-video social media platforms. Our aim was to provide the community with a dataset with well-defined yet non-standard human actions that complement common actions already represented in current action recognition datasets, thus adding variety to the actions that action recognition models are currently evaluated for.
2.2 Computer Vision Foundation Models
Foundation models have become essential tools in the field of computer vision. The advent of models like Florence[10] showcases the versatility and scope of these models. Specifically, Florence encapsulates the progression of foundation models, adeptly offering transitions from broad scene outlines to detailed object-centric views and from static imagery to dynamic video sequences. Additionally, its effectiveness spans various modalities, from RGB imagery to textual data, establishing new standards in tasks like classification, object detection, retrieval, and particularly in action recognition.
Another significant model in this domain is VideoMAE V2[9]. Introduced at CVPR 2023, this self-supervised pre-training methodology stands out for its ability to train on datasets exceeding 1 million videos. With its innovative dual masking strategy and progressive training approach, VideoMAE V2 sets new standards in action recognition. Empirical evaluations demonstrate its superior performance across multiple benchmarks, emphasizing its effectiveness in real-world scenarios. We anticipate that our extensive TikTok dataset will serve as a valuable resource for the continued development and refinement of computer vision foundation models, contributing to the advancement of this rapidly evolving field.
3 Dataset Construction
3.1 Dataset Overview
We introduce the TikTokActions dataset, designed for human action recognition. Derived from TikTok videos, this dataset comprises over 280,000 video clips from 386 hashtags that capture a wide spectrum of human actions in various scenarios. The dataset spans diverse categories, from dance movements and fitness exercises to hand gestures and daily activities. The aim of TikTokActions is to provide a robust and comprehensive foundation for training and evaluating human action recognition models in realistic settings.
3.2 How the Dataset was Built
To identify a list of hashtags to mine on TikTok, we initially consulted several commonly used datasets related to human actions, including UCF101[7], HMDB51[5], and Google DeepMind Kinetics[4]. After conducting a thorough comparison of these datasets, we identified four categories of videos that can best represent various human actions: dance, sports, fitness, and kinetics. Following these categorizations, we proceeded to identify several hashtags under each category. During the process of structuring our hashtag list, we drew insights from the hashtags commonly employed by TikTok users on the platform. This comparative analysis yielded six fundamental principles that guided our approach to selecting hashtags to mine for data: generalization, specification, noun substitution, noun-verb order, hashtags consolidation, and social/cultural consideration. The last principle enabled our team to include the fifth category unique to social media, encompassing a wide variety of challenge types of actions originating from distinct social and cultural contexts. Additionally, beyond these six principles, we placed significant emphasis on the number of views under each hashtag, recognizing that popularity and engagement levels are crucial indicators of the relevance and impact of the content for each hashtag. We set a 5 million view minimum threshold for hashtag inclusion, reflecting a focus on content demonstrating substantial public engagement and visibility, indicative of broad influence.
3.2.1 Generalization/Using the umbrella terms
When comparing the hashtags we generated with those used by TikTok users, we observed that hashtags related to specific actions can often be represented in a very broad or inclusive manner on the platform. For instance, the hashtag #basketball can serve as an umbrella term encompassing a wide array of hashtags, such as #playingbasketball, #dunking, #dribbling.
3.2.2 Specification
There are situations where specificity is crucial. For instance, within the broader classification of Ballroom Dance, distinct dance styles such as Waltz, Quickstep, and Tango exist. Instead of solely using the hashtag #ballroomdance, we include specific hashtags, including #chacha, #salsa, #tango, and #waltz, to ensure our datasets encompass pertinent human actions inherent to distinct dance styles.
3.2.3 Noun Substitution
While gathering short TikTok videos, we observed that a more effective approach for identifying the desired content is to employ a single noun instead of an entire verb phrase. For instance, the use of the hashtag #piano suffices to compile a more extensive collection(see Table  1) of short videos instead of using #pianoplaying.
Hashtags Number of Views
#piano 44.4 Billion
#playingpiano 29.4 Million
Table 1: Table of Hashtags and Their Corresponding Number of Views
Figure 1: Flowchart depicting the hashtag selection and modification process.
3.2.4 Noun-Verb Order
We noticed a preference among users for using noun phrases over verb phrases when describing their activities. For example, TikTok users frequently use hashtags like #shoerepair instead of verb phrases like #repairingshoes.
3.2.5 Hashtag Consolidation
When composing the hashtag list, we streamlined similar hashtags into a single one if they denote the same type of human action, prioritizing those with higher views on TikTok. For example, we retain #cuttingcakes and remove similar hashtags like #cutfruit, #applecutting, and #breadcutting.
3.2.6 Social/cultural Considerations
Culturally and socially relevant hashtags were included when crafting the hashtag list. For instance, the ”flower challenge” originated as a dance challenge on TikTok initiated by a K-pop star, and TikTok users subsequently adopted the hashtag #flowerchallenge when sharing their dance videos.
In Figure 1, we show an example flowchart of the hashtag selection and modification process. Our final and complete list of hashtags along with video numbers is in Appendix A.
3.3 Data Collection and Processing
We constructed the TikTokActions dataset using the official TikTok research API. For each hashtag, approximately 900 videos and corresponding metadata were mined. Each video underwent processing with tools like PySceneDetect and YOLO v8[6] producing a JSON file for every video, thereby accumulating metadata for 340,000 videos. The metadata encompassed various attributes, including the title, duration, and scene distribution based on PySceneDetect. We describe details about the preprocessing procedure below.
3.3.1 PySceneDetect and YOLOv8 Preprocessing
Our pre-processing approach began with the use of PySceneDetect to analyze the scene composition of each video. PySceneDetect was employed to identify distinct scenes within the videos, enabling us to focus on the longest scene clips. The rationale behind selecting the longest scenes was our hypothesis that these segments have the highest probability of containing human presence.
Once the longest scenes were extracted, we then utilized the YOLOv8[6] model to detect human figures in these video segments. The primary goal of this method was to accurately identify videos containing human subjects. The procedure comprised several distinct phases, each fulfilling a specific aspect of the preprocessing objective.
In the frame extraction phase, the total frame count of each video clip was determined using the ffprobe tool. Given the considerable frame count in many videos, it was computationally impractical to process each frame. Consequently, we resorted to subsampling. Typically, a fixed number of frames, defaulting to 10, were extracted at regular intervals. For a video containing 101 frames, every 10th frame was extracted starting from the 10th frame.
The next step was human detection. Here, the extracted frames underwent inference using the YOLOv8[6] model, initialized based on Ultralytics’ YOLO framework. The primary task was to review the detections and identify frames with human presence. Frames with at least one ”person” label contributed to the cumulative count of frames with human subjects for that particular video.
Lastly, the post-processing steps included thresholding. For a video to qualify as containing humans, it was required to surpass a pre-defined frame threshold with human detections, set to 4 by default. Videos below this threshold were earmarked for subsequent human scrutiny. Following the analysis of all videos, a results compilation phase calculated aggregated statistics, providing a clear picture of the number of videos with humans. Additionally, any extracted frames were systematically removed post-inference to maintain a clean working environment. In Figure  2, we show the process of the generation of clips for each video.
Figure 2: Flowchart depicting the clip extraction process from raw videos mined on TikTok.
4 Performance on Benchmarks
In the domain of human action recognition, benchmarking performance is crucial to understanding the efficacy of a model. In this section, we evaluate the performance of our model, which was initially pre-trained on VideoMAE V2[9] with TikTokActions subset and subsequently fine-tuned on two datasets: UCF-101[7] and HMDB51[5].
4.1 Pre-training on VideoMAE V2
VideoMAE V2[9] represents a state-of-the-art masked autoencoding framework with a billion-level ViT backbone. This framework has been instrumental in enhancing performance across a diverse range of video-related tasks, including video action recognition, action detection, and temporal action detection. The generalization ability of VideoMAE V2 pre-trained ViTs[1] as video foundation models has been verified across diverse downstream tasks, showcasing its robustness and adaptability.
Given the strength of VideoMAE V2[9], we utilized it for our pre-training phase. Table 2 displays the hyperparameters used for this phase. The weights of VideoMAE V2 were randomly initialized.
Table 2: Hyperparameters for Pre-training on VideoMAE V2
Parameter Value
Mask Ratio 0.9
Decoder Mask Ratio 0.5
Model Vit_base
Decoder Depth 4
Batch Size 32
Number of Frames 16
Sampling Rate 4
Number of Samples 2
Number of Workers 24
Optimizer adamw
Learning Rate 1e-3
Optimizer Betas 0.9, 0.95
Warmup Epochs 20
Total Epochs 200
We conducted our training on a single NVIDIA A100 40GB GPU. This constraint necessitated a reduction in the scale of our dataset; consequently, we focused on a randomly selected subset comprising 10,000 videos, selected from the total collection of 280,000. This approach allowed us to manage the computational demands while still providing a representative sample for training and evaluating the model’s performance. This constraint also enabled us to study the utility of self-supervised pre-training on smaller datasets and compare it to prior work which are pre-trained on larger datasets.
4.2 Fine-Tuning on Established Datasets
Parameter Value
Batch Size 8
Input Size 224
Short Side Size 224
Number of Frames 16
Sampling Rate 4
Number of Workers 24
Optimizer adamw
Learning Rate 1e-3
Drop Path 0.35
Layer Decay 0.9
Optimizer Betas 0.9, 0.999
Weight Decay 0.05
Total Epochs 100
Test Number of Segments 5
Test Number of Crops 3
Table 3: Hyperparameters for Fine-tuning
Dataset Top-1 (%) Top-5 (%)
HMDB-51[5] 53.24 89.73
UCF-101[7] 95.3 99.54
Table 4: Fine-tuning Results With 10,000 Tiktok Actions videos Pre-training
After the initial pre-training on VideoMAE V2[9], the model was fine-tuned on established datasets. The hyperparameters for the fine-tuning phase are provided in Table 3. The results of the fine-tuning on various datasets, UCF-101[7] and HMDB51[5] are consolidated in Table 4. It showcases the effectiveness of the TikTok Actions dataset in model pre-training. Notably, the model achieved a high Top-1 accuracy of 95.3% on UCF-101 and 53.24% on HMDB-51, demonstrating its strong generalization capabilities across diverse action recognition tasks.
5 Extrapolating the Required Pre-Training Dataset Size
In Figure LABEL:fig:accplot, we find a monotone increasing accuracy score versus data volume. In the spirit of Floares et al. [2], we attempt to extrapolate the number of preprocessed videos for a desired accuracy by curve fitting to the number of data points trained so far. Because the accuracy maxes out at 100%, we choose to fit curves that are monotonically increasing with a horizontal asymptote at $11$. This makes any CDF of a random variable a viable choice for curve fitting. By iterating over the random variables in Scipy [8], we choose the top 10 distributions that fit the accuracy curve. Table 5 displays the number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
We validated these projections by pre-training a model with 9157 data points and fine-tuning the model using the same methods applied in the initial projections. The outcome of this validation was notably consistent with our predictive model. The model achieved an accuracy of 94.93%, closely aligning with the expected performance as indicated by the inverse CDF projections for this dataset size.
This validation not only reinforces the reliability of our statistical model but also highlights the practicality of using such predictive techniques in determining dataset size requirements. The fact that a dataset size of just 9157 points could yield an accuracy so close to 95% – a benchmark often targeted in action recognition tasks – is a testament to the efficiency of our approach. These results suggest that the requirement for massive datasets in the pre-training phase may be reconsidered, especially when computational resources are limited. This could mark a significant shift in how datasets are compiled and utilized in machine learning, particularly in fields where data acquisition is a challenge.
Table 5: Number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
Model Desired Accuracy
95% 96% 97% 98% 99%
t
9157 12388 18323 31892 82574
Burr XII
9090 12209 17857 30511 76240
InvGamma
9072 12166 17750 30204 74859
BetaPrime
9072 12165 17748 30200 74843
F
9072 12166 17749 30203 74855
GenPareto
9072 12164 17744 30184 74759
Lomax
9072 12164 17744 30184 74759
LogLaplace
9072 12164 17744 30184 74760
Pareto
9072 12164 17743 30183 74755
InvWeibull
9043 12101 17596 29772 72915
6 Discussion
The development of the TikTok-sourced dataset aimed to contribute to the action recognition field by including a variety of well-defined but non-standard human actions. These actions were selectively curated to complement the actions typically found in existing datasets, rather than duplicating them. This approach in dataset composition was chosen to diversify the range of human actions available for training and evaluating action recognition models.
By integrating these unique and diverse actions from the TikTok platform, our dataset broadens the scope of actions recognizable by machine learning models. This expansion is not merely a quantitative increase in data but a qualitative enrichment that challenges and enhances the adaptability and comprehensiveness of action recognition algorithms. This contribution is particularly significant in light of the evolving nature of human actions in digital media, where new and unconventional actions frequently emerge.
We also made auxiliary findings with notable implications: we observed a diminished necessity for large pre-training dataset sizes. Even with a relatively smaller subset of the TikTokActions dataset, the model performance remains competitive. This challenges the prevailing notion in action recognition research that larger datasets are invariably superior for pre-training. The implications of this finding indicate a potential shift in the paradigm of dataset compilation and model training. It points towards the efficiency of a smaller, well-curated dataset that can significantly reduce computational resources while maintaining high accuracy, which may be preferred over a larger but less valuable dataset.
One of the primary limitations of this study is the incomplete pre-training on the full TikTokActions dataset. Due to computational constraints, our experiments were conducted on a subset of 10,000 videos, which might not fully represent the diversity and complexity of the full dataset. Additionally, while the dataset performed well on benchmarks like UCF101[9] and HMDB51[5], its efficacy across other popular action recognition datasets remains untested.
Another significant limitation is the absence of ground truth labels for each clip in the TikTokActions dataset. Although the videos were curated scientifically, the lack of verified labels poses a challenge for more rigorous supervised learning approaches, potentially impacting the precision of the training and evaluation processes.
Future research endeavors can utilize the entirety of or subsets of the TikTokActions dataset for pre-training, aiming to enhance model performance and generalization across action recognition tasks. Another opportunity is the integration of TikTokActions with other datasets such as Kinetics[4], which will provide insights into the unique features and potential synergies of these datasets. This comparison is crucial for validating TikTokActions’ effectiveness and identifying areas for expansion or enhancement.
An emerging area of exploration is the application of weekly self-supervised learning techniques. These methods, which involve regular, unsupervised updates to the model based on new data influxes, could significantly enhance the adaptability and accuracy of the models trained on TikTokActions.
Furthermore, employing crowd-sourcing to acquire ground truth labels for the TikTokActions dataset presents a promising opportunity. This initiative would enable researchers to conduct more supervised learning studies, potentially leading to more accurate and reliable models. By combining crowd-sourced annotations with sophisticated learning algorithms, the field can expect to see substantial advancements in action recognition capabilities and applications.
Appendix A Appendix
In the appendix, we will present additional insights into our dataset, specifically focusing on visual and statistical aspects. This includes select screenshots of our video clips, which exemplify the diversity and context of the actions captured within our dataset. Alongside these visuals, we will include the figure, which provides the number distributions of videos for each hashtag. Furthermore, we will present the distribution of video durations, offering a comprehensive view of the temporal characteristics of our clips. These elements combined will give a more complete picture of the dataset’s richness.
Appendix B Sample Clips
In Figure 3, we present the extensive diversity of the TikTokActions dataset, which captures a wide range of human activities tagged across multiple domains such as fitness, sports, and dance. The collection features a variety of video types that include full-body human actions, clips focusing on hand movements, and scenarios involving multiple individuals. The presence of such varied content provides a rich learning ground for self-supervised learning models to extract and learn a comprehensive set of features from human actions, gestures, and interactions, aiding in the enhancement of action recognition capabilities.
Figure 3: Example classes from the TikTokActions dataset.
Appendix C Dataset Composition
Our dataset offers a diverse array of video counts across different hashtags. Figure 4 illustrates the distribution of videos across various hashtags, indicating the prevalence of certain actions within our dataset. On average, each hashtag is associated with approximately 735 videos. The ’sax’ category stands out with the maximum count of 938 videos, while the ’mopping’ category has the minimum, with 325 videos.
Figure 4: Bar chart showing the number of videos per hashtag in our dataset, demonstrating the frequency and distribution of action categories.
Conversely, Figure 5 displays a pie chart that categorizes the dataset into four video count ranges: 200-400, 400-600, 600-800, and 800+ videos per category. It is noteworthy that a significant portion, 53.4%, of the hashtags have a video count that falls within the 600-800 range.
Figure 5: Pie chart categorization of video counts in different ranges, highlighting the dataset’s variety and the volume of data available for model training.
Appendix D Dataset Duration
In the composition of the TikTokActions dataset, a key focus was on ensuring that each video clip is sufficiently long to capture a complete action, while also maintaining a manageable length for analysis. To this end, we established a minimum clip duration of 3.5 seconds, ensuring each clip is long enough to encompass a full action. Simultaneously, to avoid excessively long clips, which can complicate analysis and may contain extraneous information, we capped the maximum duration at 10 seconds. For videos exceeding this length, only the most relevant 10-second segment, typically featuring the most significant action, was selected. This methodology results in a dataset that is both comprehensive and tailored for efficient action recognition analysis.
The following table details the distribution of video durations within our dataset, post these adjustments:
Duration Range (seconds) Number of Videos
3.5-5 30103
5-10 81356
10+ 172123
Table 6: Distribution of video durations in the TikTokActions dataset. The data reflects the adherence to the minimum and maximum duration thresholds for clip selection.

### v2

@@
 By combining crowd-sourced annotations with sophisticated learning algorithms, the field can expect to see substantial advancements in action recognition capabilities and applications.
+Acknowledgements
+This work was made possible by a multitude of supporters and funding sources.
+We thank the National Science Foundation for their generous support through the CyberTraining grant (Award No.
+2118222) under the HI-DSI project.
+The technical assistance and advanced computing resources provided by the University of Hawaii Information Technology Services – Cyberinfrastructure, partially funded by NSF CC* awards #2201428 and #2232862 (Koa), have been invaluable.
+The development and acceleration of cloud computing through JetStream2 have significantly contributed to our research; we acknowledge David Y.
+Hancock, Jeremy Fischer, John Michael Lowe, Winona Snapp-Childs, Marlon Pierce, Suresh Marru, J.
+Eric Coulter, Matthew Vaughn, Brian Beck, Nirav Merchant, Edwin Skidmore, and Gwen Jacobs for this computational infrastructure (DOI: https://doi.org/10.1145/3437359.3465565).
+The AWS Cloud Credit for Research program has also played a crucial role in providing computational resources for our work.
+We are grateful for the funding provided by the National Institutes of Health under the NIH Director’s New Innovator Award (award DP2-EB035858).
 Appendix A Appendix

### v3

@@
-TikTokActions: A TikTok-Derived Video Dataset for Human Action Recognition
+Advancing Human Action Recognition with Foundation Models trained on Unlabeled Public Videos
 Abstract
-The increasing variety and quantity of tagged multimedia content on platforms such as TikTok provides an opportunity to advance computer vision modeling.
-We have curated a distinctive dataset of 283,582 unique video clips categorized under 386 hashtags relating to modern human actions.
-We release this dataset as a valuable resource for building domain-specific foundation models for human movement modeling tasks such as action recognition.
-To validate this dataset, which we name TikTokActions, we perform two sets of experiments.
-First, we pretrain the state-of-the-art VideoMAEv2 with a ViT-base backbone on TikTokActions subset, and then fine-tune and evaluate on popular datasets such as UCF101 and the HMDB51.
-We find that the performance of the model pre-trained using our TikTok dataset is comparable to models trained on larger action recognition datasets(95.3% on UCF101 and 53.24% on HMDB51).
-Furthermore, our investigation into the relationship between pre-training dataset size and fine-tuning performance reveals that beyond a certain threshold, the incremental benefit of larger training sets diminishes.
-This work introduces a useful TikTok video dataset that is available for public use and provides insights into the marginal benefit of increasing pre-training dataset sizes for video-based foundation models.
+The increasing variety and quantity of tagged multimedia content on a variety of online platforms offers a unique opportunity to advance the field of human action recognition.
+In this study, we utilize 283,582 unique, unlabeled TikTok video clips, categorized into 386 hashtags, to train a domain-specific foundation model for action recognition.
+We employ VideoMAEv2, an advanced model integrating Masked Autoencoders (MAE) with Vision Transformers (ViT), pre-trained on this diverse collection of unstructured videos.
+Our model, fine-tuned on established action recognition benchmarks such as UCF101 and HMDB51, achieves results comparable with the state-of-the-art: 97.5% on UCF101, 83.9% on HMDB51, 81.1% on Kinetics-400, and 69.2% on Something-Something V2 using the ViT-base backbone.
+These results highlight the potential of using TikTok videos as a valuable source of diverse and dynamic content for training foundation models.
+Our investigation confirms that while initial increases in pre-training data volume significantly enhance model performance, the gains diminish as the dataset size continues to expand.
+Our findings identify two critical lessons in self-supervised learning for computer vision: (1) additional pre-training data yields diminishing benefits and (2) these diminishing benefits can be observed at relatively small dataset sizes on the order of thousands of images.
 1 Introduction
-Action recognition is a fundamental computer vision task that involves identifying and classifying human actions in video sequences, with significant applications in areas such as surveillance, interactive gaming, and healthcare.
-In the era of social media platforms, TikTok has emerged as a significant repository of human action videos.
-We have compiled a dataset of 283,582 TikTok video clips that encompass a range of human activities, including dance, sports, fitness, hand gestures, and daily routines.
-This collection aims to represent various human actions in a manner that is relevant to real-world scenarios and cultural contexts.
-Importantly, this dataset is uniquely curated to include well-defined but non-standard human actions, filling gaps in existing action recognition datasets and diversifying the range of human activities recognized by current models.
-We developed a model for human action recognition trained using the curated TikTok video clips.
-Pre-training on a dataset with varied and authentic content is intended to more accurately reflect the spectrum of naturalistic human actions, aiming to address the limitations of conventional datasets by incorporating elements of contemporary behavior.
-In our initial evaluations, presented later in this paper, we find that utilizing our TikTok dataset may enhance the development of targeted human action recognition models in the field of computer vision.
-Moreover, the nature of our data source, TikTok, presents another advantage.
-Kinetics[4] derives its content from YouTube, featuring a combination of professional and amateur videos, whereas TikTok’s database consists primarily of user-generated material.
-This means that our dataset is inherently rich with real-world scenarios: from spontaneous camera movements and varied illumination conditions to diverse backgrounds and authentic human interactions.
-This rawness provides a genuine challenge and opportunity for training robust models capable of understanding human actions in a plethora of real-world conditions.
-We conducted a series of experiments to assess the quality of our dataset and its potential to serve as a basis for pre-training foundation models for action-related tasks.
-We used a well-established state-of-the-art model VideoMAE V2[9], and pre-trained it on our TikTokActions dataset.
-We then fine-tuned the model on two well-known video action recognition datasets: UCF101[7] and HMDB51[5].
-Our experiments provided valuable insights: the performance achieved was comparable to models that were both pre-trained and fine-tuned solely on UCF101 and HMDB51.
-This underscores the utility of the TikTokActions dataset to serve as a useful pre-training dataset for action-related tasks.
-We next delved into understanding the relationship between the quantity of pre-training videos and the performance on the downstream action recognition task.
-By sampling subsets of our TikTok dataset, starting with 1,000 videos and increasing in increments of 1,000 up to 6,000 videos, pre-training VideoMAE V2[9] on the sampled subset, and subsequently fine-tuning each on the benchmark datasets, we observed that as long as the number of pre-training videos remains above a certain threshold, the performance remains comparable to massive pre-training datasets.
-This provides a valuable lesson when pre-training foundation models: the benefits of increasing the pre-training dataset size diminish as the dataset increases.
+Action recognition is a fundamental computer vision task that involves identifying and classifying human actions in video sequences, with significant applications in areas such as security, interactive gaming, and healthcare.
+In the era of social media platforms, TikTok has emerged as a significant source of human action videos.
+We utilized a collection of 283,582 TikTok video clips, encompassing a wide range of human activities, to train a foundation model tailored to human action recognition.
+This internally curated collection aims to represent various human actions relevant to real-world scenarios and cultural contexts.
+Importantly, it includes well-defined but non-standard human actions, addressing gaps in existing action recognition datasets and diversifying the range of human activities recognized by current models.
+Unlike Kinetics [4], which also includes user-generated content from YouTube, our collection exclusively features TikTok videos, known for their dynamic and spontaneous nature.
+This distinction offers unique training advantages, as such videos often exhibit quick transitions, novel perspectives, and creative interactions—elements less prevalent in YouTube’s often longer-form content.
+These characteristics provide robust challenges for training models, enhancing their ability to interpret a wide range of human actions in diverse and unpredictable environments.
+We conducted a series of experiments to assess the potential of videos collected in this manner for pre-training computer vision foundation models for action-related tasks.
+We employed the well-established, state-of-the-art model VideoMAE V2, using it with ViT-base backbones[9].
+The ViT-base was used to provide a baseline comparison with other models using similarly sized backbones.
+Subsequent fine-tuning on well-known video action recognition datasets like UCF101[7], HMDB51[5], Kinetics[4], and Something-Something V2[3] demonstrated valuable insights: the performance achieved with both backbones was comparable to, or in some cases even better than, models trained exclusively on these conventional datasets or even those pre-trained on much larger datasets, where performance increases are often marginal.
+We also explored the relationship between the quantity of pre-training videos and the performance on downstream tasks.
+By sampling subsets of our collection, starting with 1,000 videos and increasing in increments of 1,000 up to 6,000 videos, and pre-training VideoMAE V2[9] on these subsets, we observed that performance remains robust as long as the number of pre-training videos stays above a certain threshold, highlighting that massive datasets may not always be necessary for effective pre-training.
+This exploration not only emphasizes the efficiency of using online unlabeled videos as a pre-training source but also challenges the conventional reliance on extensive video datasets for achieving high model performance.
+Before deployment of this approach in the real world, we encourage researchers to think through the ethics of data scraping from publicly available sources to train deep learning models.
 2 Related Work
@@
 3 Dataset Construction
-3.1 Dataset Overview
-We introduce the TikTokActions dataset, designed for human action recognition.
-Derived from TikTok videos, this dataset comprises over 280,000 video clips from 386 hashtags that capture a wide spectrum of human actions in various scenarios.
-The dataset spans diverse categories, from dance movements and fitness exercises to hand gestures and daily activities.
-The aim of TikTokActions is to provide a robust and comprehensive foundation for training and evaluating human action recognition models in realistic settings.
-3.2 How the Dataset was Built
-To identify a list of hashtags to mine on TikTok, we initially consulted several commonly used datasets related to human actions, including UCF101[7], HMDB51[5], and Google DeepMind Kinetics[4].
-After conducting a thorough comparison of these datasets, we identified four categories of videos that can best represent various human actions: dance, sports, fitness, and kinetics.
-Following these categorizations, we proceeded to identify several hashtags under each category.
-During the process of structuring our hashtag list, we drew insights from the hashtags commonly employed by TikTok users on the platform.
-This comparative analysis yielded six fundamental principles that guided our approach to selecting hashtags to mine for data: generalization, specification, noun substitution, noun-verb order, hashtags consolidation, and social/cultural consideration.
-The last principle enabled our team to include the fifth category unique to social media, encompassing a wide variety of challenge types of actions originating from distinct social and cultural contexts.
-Additionally, beyond these six principles, we placed significant emphasis on the number of views under each hashtag, recognizing that popularity and engagement levels are crucial indicators of the relevance and impact of the content for each hashtag.
-We set a 5 million view minimum threshold for hashtag inclusion, reflecting a focus on content demonstrating substantial public engagement and visibility, indicative of broad influence.
+3.1 Video Collection Overview
+To support our experimentation exploring the utility of unlabeled public data for pre-training foundation models for human action recognition, we curated an extensive collection of TikTok videos.
+This collection, consisting of over 280,000 video clips from 386 hashtags, captures a wide spectrum of human actions across various scenarios.
+The videos span diverse categories, from dance movements and fitness exercises to hand gestures and daily activities.
+Unlike traditional datasets like Kinetics, our collection uniquely incorporates elements of pop culture, memes, and rapid trend evolution, providing dynamic content that significantly boosts the model’s ability to understand and generalize from real-world, culturally relevant scenarios.
+3.2 Video Selection Methodology
+To curate a diverse and relevant set of videos from TikTok, we initially examined several well-known action recognition datasets, including UCF101[7], HMDB51[5], and Google DeepMind Kinetics[4].
+These datasets helped us to understand the broad categories of human actions typically represented in research.
+We identified four primary categories: dance, sports, fitness, and kinetics.
+Building upon these categories, we explored TikTok to identify relevant hashtags that capture a wide array of human actions within these domains.
+Our selection was strategically guided by six principles designed to ensure a comprehensive and culturally relevant collection: generalization, specification, noun substitution, noun-verb order, hashtag consolidation, and social/cultural relevance.
+These principles helped ensure that the videos not only cover a broad spectrum of actions but also include dynamic and culturally significant content such as memes, pop culture trends, and unique challenges—elements that are distinctly prevalent on TikTok.
+The social/cultural relevance was particularly important, allowing us to incorporate an additional category that includes a wide variety of actions originating from distinct social and cultural contexts, such as viral dance challenges and fitness trends.
+Popularity and engagement levels also played a crucial role in our selection process; we set a minimum threshold of 5 million views per hashtag to ensure that the content is not only relevant but also resonates widely with users, reflecting substantial public engagement and visibility.
+This approach allows us to harness rich, culturally nuanced content that can aid in training models to better understand and interpret human actions in real-world contexts.
 3.2.1 Generalization/Using the umbrella terms
@@
 #playingpiano 29.4 Million
-Table 1: Table of Hashtags and Their Corresponding Number of Views
-Figure 1: Flowchart depicting the hashtag selection and modification process.
+Table 1: Example of the expansion in data availability when removing nouns from the hashtag construction process.
+Figure 1: The hashtag selection and modification process.
 3.2.4 Noun-Verb Order
@@
 In Figure 1, we show an example flowchart of the hashtag selection and modification process. Our final and complete list of hashtags along with video numbers is in Appendix A.
-3.3 Data Collection and Processing
-We constructed the TikTokActions dataset using the official TikTok research API.
-For each hashtag, approximately 900 videos and corresponding metadata were mined.
-Each video underwent processing with tools like PySceneDetect and YOLO v8[6] producing a JSON file for every video, thereby accumulating metadata for 340,000 videos.
-The metadata encompassed various attributes, including the title, duration, and scene distribution based on PySceneDetect.
-We describe details about the preprocessing procedure below.
-3.3.1 PySceneDetect and YOLOv8 Preprocessing
-Our pre-processing approach began with the use of PySceneDetect to analyze the scene composition of each video.
-PySceneDetect was employed to identify distinct scenes within the videos, enabling us to focus on the longest scene clips.
-The rationale behind selecting the longest scenes was our hypothesis that these segments have the highest probability of containing human presence.
-Once the longest scenes were extracted, we then utilized the YOLOv8[6] model to detect human figures in these video segments.
-The primary goal of this method was to accurately identify videos containing human subjects.
-The procedure comprised several distinct phases, each fulfilling a specific aspect of the preprocessing objective.
-In the frame extraction phase, the total frame count of each video clip was determined using the ffprobe tool.
-Given the considerable frame count in many videos, it was computationally impractical to process each frame.
-Consequently, we resorted to subsampling.
-Typically, a fixed number of frames, defaulting to 10, were extracted at regular intervals.
-For a video containing 101 frames, every 10th frame was extracted starting from the 10th frame.
-The next step was human detection.
-Here, the extracted frames underwent inference using the YOLOv8[6] model, initialized based on Ultralytics’ YOLO framework.
-The primary task was to review the detections and identify frames with human presence.
-Frames with at least one ”person” label contributed to the cumulative count of frames with human subjects for that particular video.
-Lastly, the post-processing steps included thresholding.
-For a video to qualify as containing humans, it was required to surpass a pre-defined frame threshold with human detections, set to 4 by default.
-Videos below this threshold were earmarked for subsequent human scrutiny.
-Following the analysis of all videos, a results compilation phase calculated aggregated statistics, providing a clear picture of the number of videos with humans.
-Additionally, any extracted frames were systematically removed post-inference to maintain a clean working environment.
-In Figure  2, we show the process of the generation of clips for each video.
-Figure 2: Flowchart depicting the clip extraction process from raw videos mined on TikTok.
+3.3 Video Collection and Processing
+We curated a comprehensive video collection using the official TikTok research API, focusing on approximately 900 videos per hashtag from a set of hashtags identified through our selection process.
+Each video and its metadata were processed to facilitate model training, using tools such as PySceneDetect and YOLO v8[6].
+This processing resulted in a JSON file for each video, which included metadata capturing attributes like title, duration, and scene distribution, as analyzed by PySceneDetect.
+3.3.1 PySceneDetect and YOLOv8 Processing
+Our processing approach began with PySceneDetect, used to analyze the scene composition of each video.
+This tool helped identify distinct scenes within the videos, enabling us to focus on the longest scene clips, hypothesized to have the highest likelihood of containing significant human action.
+Once the longest scenes were identified, the YOLOv8[6] model was employed to detect human figures in these segments.
+The primary goal of this processing was to ensure accurate identification of videos containing human subjects, vital for the training of our models.
+During the frame extraction phase, we used the ffprobe tool to determine the total frame count of each video clip.
+Due to the large number of frames in many videos, we applied subsampling for efficiency.
+Typically, a fixed number of frames (usually 10) were extracted at regular intervals throughout the video.
+The extracted frames were then processed using the YOLOv8[6] model to detect human presence.
+Frames that included at least one ”person” label were counted towards the total number of frames containing humans for that particular video.
+Post-processing included applying a threshold for human presence; videos needed to exceed a pre-defined frame threshold with human detections to qualify.
+Videos that did not meet this threshold were subject to further review.
+This systematic approach helped compile aggregated statistics on the human presence in our video collection, ensuring the quality of content used for model training.
+The entire process, from video selection to frame extraction and human detection, was documented to maintain transparency and reproducibility.
+Figure  2 illustrates the process of clip extraction from the raw videos sourced from TikTok.
+Figure 2: The clip extraction process from raw videos mined on TikTok.
 4 Performance on Benchmarks
-In the domain of human action recognition, benchmarking performance is crucial to understanding the efficacy of a model.
-In this section, we evaluate the performance of our model, which was initially pre-trained on VideoMAE V2[9] with TikTokActions subset and subsequently fine-tuned on two datasets: UCF-101[7] and HMDB51[5].
+In the domain of human action recognition, benchmarking performance is essential to understand the efficacy of a model.
+This section evaluates the performance of our model, initially pre-trained on the entire collection of 283,582 unlabeled TikTok video clips using VideoMAE V2[9], and subsequently fine-tuned on Kinetics-400 (K400).
+Following this, the model underwent further fine-tuning on UCF-101[7] and HMDB51[5], adopting a transfer learning approach akin to the experimental setups used in VideoMAE V2 studies.
+Additionally, we extended our evaluation to include the Something-Something V2[3] dataset, assessing the model’s adaptability and performance across varied action recognition contexts.
 4.1 Pre-training on VideoMAE V2
-VideoMAE V2[9] represents a state-of-the-art masked autoencoding framework with a billion-level ViT backbone.
-This framework has been instrumental in enhancing performance across a diverse range of video-related tasks, including video action recognition, action detection, and temporal action detection.
-The generalization ability of VideoMAE V2 pre-trained ViTs[1] as video foundation models has been verified across diverse downstream tasks, showcasing its robustness and adaptability.
-Given the strength of VideoMAE V2[9], we utilized it for our pre-training phase. Table 2 displays the hyperparameters used for this phase. The weights of VideoMAE V2 were randomly initialized.
-Table 2: Hyperparameters for Pre-training on VideoMAE V2
-Parameter Value
-Mask Ratio 0.9
-Decoder Mask Ratio 0.5
-Model Vit_base
-Decoder Depth 4
-Batch Size 32
-Number of Frames 16
-Sampling Rate 4
-Number of Samples 2
-Number of Workers 24
-Optimizer adamw
-Learning Rate 1e-3
-Optimizer Betas 0.9, 0.95
-Warmup Epochs 20
-Total Epochs 200
-We conducted our training on a single NVIDIA A100 40GB GPU.
-This constraint necessitated a reduction in the scale of our dataset; consequently, we focused on a randomly selected subset comprising 10,000 videos, selected from the total collection of 280,000.
-This approach allowed us to manage the computational demands while still providing a representative sample for training and evaluating the model’s performance.
-This constraint also enabled us to study the utility of self-supervised pre-training on smaller datasets and compare it to prior work which are pre-trained on larger datasets.
+VideoMAE V2[9] is a state-of-the-art masked autoencoding framework, notable for its flexibility in backbone selection, ranging from base to billion-level configurations.
+We opted for the ViT-base backbone, which, while less computationally intense than its billion-level counterparts, significantly enhances performance across various video-related tasks, including video action recognition.
+The generalization capability of VideoMAE V2 pre-trained ViTs[1] as video foundation models has been demonstrated across multiple downstream tasks, highlighting its robustness and adaptability.
+Our training was conducted on 8 NVIDIA A100 40GB GPUs, utilizing the full set of 283,582 unlabeled TikTok videos.
+The pre-training on the ViT-base backbone lasted for 20 days, providing a comprehensive base for developing a robust model capable of understanding a wide array of human actions.
 4.2 Fine-Tuning on Established Datasets
-Parameter Value
-Batch Size 8
-Input Size 224
-Short Side Size 224
-Number of Frames 16
-Sampling Rate 4
-Number of Workers 24
-Optimizer adamw
-Learning Rate 1e-3
-Drop Path 0.35
-Layer Decay 0.9
-Optimizer Betas 0.9, 0.999
-Weight Decay 0.05
-Total Epochs 100
-Test Number of Segments 5
-Test Number of Crops 3
-Table 3: Hyperparameters for Fine-tuning
+After initial pre-training, the model underwent two distinct paths of fine-tuning.
+The first path involved direct fine-tuning on the Kinetics-400 and Something-Something V2 datasets to assess immediate applicability to broad action recognition tasks.
+The results of this direct fine-tuning are detailed in Table 2.
+The second path involved using the model fine-tuned on Kinetics-400 as an intermediary for transfer learning on UCF-101 and HMDB51, leveraging robust pre-training on TikTok videos.
+The outcomes of this transfer learning process are presented in Table 3.
 Dataset Top-1 (%) Top-5 (%)
-HMDB-51[5] 53.24 89.73
-UCF-101[7] 95.3 99.54
-Table 4: Fine-tuning Results With 10,000 Tiktok Actions videos Pre-training
-After the initial pre-training on VideoMAE V2[9], the model was fine-tuned on established datasets.
-The hyperparameters for the fine-tuning phase are provided in Table 3.
-The results of the fine-tuning on various datasets, UCF-101[7] and HMDB51[5] are consolidated in Table 4.
-It showcases the effectiveness of the TikTok Actions dataset in model pre-training.
-Notably, the model achieved a high Top-1 accuracy of 95.3% on UCF-101 and 53.24% on HMDB-51, demonstrating its strong generalization capabilities across diverse action recognition tasks.
+Kinetics-400[4] 81.1 94.8
+Something-Something V2[3] 69.2 91.6
+Table 2: Direct Fine-Tuning Results With TikTok Videos Pre-training
+Dataset Top-1 (%) Top-5 (%)
+HMDB-51[5] 83.9 96.7
+UCF-101[7] 97.5 99.6
+Table 3: Transfer Learning Results From K400 Fine-Tuning Through TikTok Pre-training
+The fine-tuning results highlight the significant potential of using large, unlabeled video collections from platforms like TikTok for advancing human action recognition tasks.
+This approach capitalizes on the rich, varied nature of content available on social media, which, when used for pre-training, enhances the model’s ability to generalize across different datasets and recognition scenarios.
+The robust performance across both direct and transfer learning paths underscores the value of TikTok videos in training models to tackle complex action recognition challenges.
 5 Extrapolating the Required Pre-Training Dataset Size
@@
 95% 96% 97% 98% 99%
-t
-9157 12388 18323 31892 82574
-Burr XII
-9090 12209 17857 30511 76240
-InvGamma
-9072 12166 17750 30204 74859
-BetaPrime
-9072 12165 17748 30200 74843
-F
-9072 12166 17749 30203 74855
-GenPareto
-9072 12164 17744 30184 74759
-Lomax
-9072 12164 17744 30184 74759
-LogLaplace
-9072 12164 17744 30184 74760
-Pareto
-9072 12164 17743 30183 74755
-InvWeibull
-9043 12101 17596 29772 72915
+t 9157 12388 18323 31892 82574
+Burr XII 9090 12209 17857 30511 76240
+InvGamma 9072 12166 17750 30204 74859
+BetaPrime 9072 12165 17748 30200 74843
+F 9072 12166 17749 30203 74855
+GenPareto 9072 12164 17744 30184 74759
+Lomax 9072 12164 17744 30184 74759
+LogLaplace 9072 12164 17744 30184 74760
+Pareto 9072 12164 17743 30183 74755
+InvWeibull 9043 12101 17596 29772 72915
 6 Discussion
-The development of the TikTok-sourced dataset aimed to contribute to the action recognition field by including a variety of well-defined but non-standard human actions.
-These actions were selectively curated to complement the actions typically found in existing datasets, rather than duplicating them.
-This approach in dataset composition was chosen to diversify the range of human actions available for training and evaluating action recognition models.
-By integrating these unique and diverse actions from the TikTok platform, our dataset broadens the scope of actions recognizable by machine learning models.
-This expansion is not merely a quantitative increase in data but a qualitative enrichment that challenges and enhances the adaptability and comprehensiveness of action recognition algorithms.
-This contribution is particularly significant in light of the evolving nature of human actions in digital media, where new and unconventional actions frequently emerge.
-We also made auxiliary findings with notable implications: we observed a diminished necessity for large pre-training dataset sizes.
-Even with a relatively smaller subset of the TikTokActions dataset, the model performance remains competitive.
-This challenges the prevailing notion in action recognition research that larger datasets are invariably superior for pre-training.
-The implications of this finding indicate a potential shift in the paradigm of dataset compilation and model training.
-It points towards the efficiency of a smaller, well-curated dataset that can significantly reduce computational resources while maintaining high accuracy, which may be preferred over a larger but less valuable dataset.
-One of the primary limitations of this study is the incomplete pre-training on the full TikTokActions dataset.
-Due to computational constraints, our experiments were conducted on a subset of 10,000 videos, which might not fully represent the diversity and complexity of the full dataset.
-Additionally, while the dataset performed well on benchmarks like UCF101[9] and HMDB51[5], its efficacy across other popular action recognition datasets remains untested.
-Another significant limitation is the absence of ground truth labels for each clip in the TikTokActions dataset.
-Although the videos were curated scientifically, the lack of verified labels poses a challenge for more rigorous supervised learning approaches, potentially impacting the precision of the training and evaluation processes.
-Future research endeavors can utilize the entirety of or subsets of the TikTokActions dataset for pre-training, aiming to enhance model performance and generalization across action recognition tasks.
-Another opportunity is the integration of TikTokActions with other datasets such as Kinetics[4], which will provide insights into the unique features and potential synergies of these datasets.
-This comparison is crucial for validating TikTokActions’ effectiveness and identifying areas for expansion or enhancement.
-An emerging area of exploration is the application of weekly self-supervised learning techniques.
-These methods, which involve regular, unsupervised updates to the model based on new data influxes, could significantly enhance the adaptability and accuracy of the models trained on TikTokActions.
-Furthermore, employing crowd-sourcing to acquire ground truth labels for the TikTokActions dataset presents a promising opportunity.
-This initiative would enable researchers to conduct more supervised learning studies, potentially leading to more accurate and reliable models.
-By combining crowd-sourced annotations with sophisticated learning algorithms, the field can expect to see substantial advancements in action recognition capabilities and applications.
+The curation of TikTok videos for enhancing human action recognition models provides a unique contribution to the field, especially by incorporating a variety of distinct yet non-traditional human actions.
+These actions, sourced from the TikTok platform, were carefully selected to complement the more conventional actions typically found in established datasets like UCF101[7] and HMDB51[5].
+This approach not only diversifies the types of human actions available for model training but also qualitatively enriches the dataset, thereby challenging and extending the adaptability of action recognition algorithms.
+Integrating these unique and diverse actions from TikTok, we have expanded the range of actions that machine learning models can recognize, addressing the dynamic nature of human actions in digital media where new and unconventional actions frequently emerge.
+This strategy highlights the value of leveraging culturally rich and dynamic user-generated content as a training resource in an era dominated by digital media.
+Furthermore, our findings challenge the conventional wisdom that larger datasets always yield better pre-training results.
+The competitive performance of our models, initially pre-trained on a curated subset of TikTok videos and subsequently fine-tuned on Kinetics-400[4] (K400) and Something-Something V2(Sth-Sth)[3], suggests that a well-curated, smaller dataset can sometimes outperform a larger, more generic one.
+This observation encourages a shift towards more efficient dataset compilation and model training paradigms, emphasizing data quality over quantity.
+Looking ahead, exploring the potential of using the full breadth of the TikTok video collection for pre-training could further enhance model performance and generalization across various tasks.
+Additionally, integrating these videos with other datasets, such as Kinetics-400 and Something-Something, could provide deeper insights into the unique characteristics and potential synergies between different types of video content.
+An emerging area of interest is the application of weekly self-supervised learning techniques, which involve regular, unsupervised updates to the model based on new data inflows.
+Such techniques could significantly boost the adaptability and accuracy of models trained with dynamic, culturally relevant content from platforms like TikTok.
+7 Ethics
+This research adheres to strict ethical guidelines regarding the use of online video data, specifically videos collected from TikTok.
+All videos were legally accessed via the TikTok Research API, and only those videos where the user granted download permission were included.
+Our data collection process is in full compliance with the TikTok Research API Terms of Service.
+As an additional layer of security, we delete the data after training.
+We place a high priority on privacy and copyright issues.
+To protect individual privacy, we have implemented face obfuscation techniques in the video clips included in our paper.
+Moreover, in line with the terms of service, we do not distribute any TikTok user content directly.
+Instead, we focus on training a foundation model, which we only use to identify action patterns, not personal biometric features.
+Ethical Considerations in Online Video Research: The use of online video data poses unique challenges and ethical considerations, particularly regarding privacy, informed consent, and the potential risks associated with the use of such data.
+Our study emphasizes the importance of adhering to ethical principles that safeguard participant interests.
+We have made significant efforts to minimize potential harm by anonymizing the data and not sharing the scraped dataset.
+Although techniques like pixelating faces can be circumvented and are not foolproof, they represent our commitment to reducing identifiable information.
+Furthermore, we advocate for the inclusion of clearer guidelines in the terms of service on video platforms, which would allow users to explicitly exclude their videos from research use.
+This approach does not replace the need for informed consent but is a step towards enhancing ethical practices in online video research.
+We aim to foster a transparent and informed discussion about the use of online video data in computational research.
 Acknowledgements
@@
 Hancock, Jeremy Fischer, John Michael Lowe, Winona Snapp-Childs, Marlon Pierce, Suresh Marru, J.
-Eric Coulter, Matthew Vaughn, Brian Beck, Nirav Merchant, Edwin Skidmore, and Gwen Jacobs for this computational infrastructure (DOI: https://doi.org/10.1145/3437359.3465565).
+Eric Coulter, Matthew Vaughn, Brian Beck, Nirav Merchant, Edwin Skidmore, and Gwen Jacobs for this computational infrastructure (DOI: 10.1145/3437359.3465565).
 The AWS Cloud Credit for Research program has also played a crucial role in providing computational resources for our work.
 We are grateful for the funding provided by the National Institutes of Health under the NIH Director’s New Innovator Award (award DP2-EB035858).
+We used ChatGPT to edit the grammar of our manuscript and to re-phrase sentences that were originally worded unclearly.
+However, all contents in this manuscript are original ideas and analyses conducted by the authors.
 Appendix A Appendix
@@
 It is noteworthy that a significant portion, 53.4%, of the hashtags have a video count that falls within the 600-800 range.
-Figure 5: Pie chart categorization of video counts in different ranges, highlighting the dataset’s variety and the volume of data available for model training.
+Figure 5: Video counts in different ranges, highlighting the dataset’s variety and the volume of data available for model training.
 Appendix D Dataset Duration

### v4

@@
 Abstract
-The increasing variety and quantity of tagged multimedia content on a variety of online platforms offers a unique opportunity to advance the field of human action recognition.
+The increasing variety and quantity of tagged multimedia content on a variety of online platforms offer a unique opportunity to advance the field of human action recognition.
 In this study, we utilize 283,582 unique, unlabeled TikTok video clips, categorized into 386 hashtags, to train a domain-specific foundation model for action recognition.
 We employ VideoMAE V2, an advanced model integrating Masked Autoencoders (MAE) with Vision Transformers (ViT), pre-trained on this diverse collection of unstructured videos.
-Our model, fine-tuned on established action recognition benchmarks such as UCF101 and HMDB51, achieves results comparable with the state-of-the-art: 97.5% on UCF101, 83.9% on HMDB51, 81.1% on Kinetics-400, and 69.2% on Something-Something V2 using the ViT-base backbone.
-These results highlight the potential of using TikTok videos as a valuable source of diverse and dynamic content for training foundation models.
+Our model, fine-tuned on established action recognition benchmarks such as UCF101 and HMDB51, achieves state-of-the-art results: 99.05% on UCF101, 86.08% on HMDB51, 85.51% on Kinetics-400, and 74.27% on Something-Something V2 using the ViT-giant backbone.
+These results highlight the potential of using unstructured and unlabeled videos as a valuable source of diverse and dynamic content for training foundation models.
 Our investigation confirms that while initial increases in pre-training data volume significantly enhance model performance, the gains diminish as the dataset size continues to expand.
-Our findings identify two critical lessons in self-supervised learning for computer vision: (1) additional pre-training data yields diminishing benefits and (2) these diminishing benefits can be observed at relatively small dataset sizes on the order of thousands of images.
+Our findings emphasize two critical axioms in self-supervised learning for computer vision: (1) additional pre-training data can yield diminishing benefits for some datasets and (2) quality is more important than quantity in self-supervised learning, especially when building foundation models.
+keywords:
+Self-Supervised Learning, Action Recognition, Vision Transformers, TikTok Video Pre-training
 1 Introduction
-Action recognition is a fundamental computer vision task that involves identifying and classifying human actions in video sequences, with significant applications in areas such as security, interactive gaming, and healthcare.
-In the era of social media platforms, TikTok has emerged as a significant source of human action videos.
-We utilized a collection of 283,582 TikTok video clips, encompassing a wide range of human activities, to train a foundation model tailored to human action recognition.
-This internally curated collection aims to represent various human actions relevant to real-world scenarios and cultural contexts.
-Importantly, it includes well-defined but non-standard human actions, addressing gaps in existing action recognition datasets and diversifying the range of human activities recognized by current models.
-Unlike Kinetics [4], which also includes user-generated content from YouTube, our collection exclusively features TikTok videos, known for their dynamic and spontaneous nature.
-This distinction offers unique training advantages, as such videos often exhibit quick transitions, novel perspectives, and creative interactions—elements less prevalent in YouTube’s often longer-form content.
-These characteristics provide robust challenges for training models, enhancing their ability to interpret a wide range of human actions in diverse and unpredictable environments.
-We conducted a series of experiments to assess the potential of videos collected in this manner for pre-training computer vision foundation models for action-related tasks.
-We employed the well-established, state-of-the-art model VideoMAE V2, using it with ViT-base backbones[9].
-The ViT-base was used to provide a baseline comparison with other models using similarly sized backbones.
-Subsequent fine-tuning on well-known video action recognition datasets like UCF101[7], HMDB51[5], Kinetics[4], and Something-Something V2[3] demonstrated valuable insights: the performance achieved with both backbones was comparable to, or in some cases even better than, models trained exclusively on these conventional datasets or even those pre-trained on much larger datasets, where performance increases are often marginal.
-We also explored the relationship between the quantity of pre-training videos and the performance on downstream tasks.
-By sampling subsets of our collection, starting with 1,000 videos and increasing in increments of 1,000 up to 6,000 videos, and pre-training VideoMAE V2[9] on these subsets, we observed that performance remains robust as long as the number of pre-training videos stays above a certain threshold, highlighting that massive datasets may not always be necessary for effective pre-training.
-This exploration not only emphasizes the efficiency of using online unlabeled videos as a pre-training source but also challenges the conventional reliance on extensive video datasets for achieving high model performance.
-Before deployment of this approach in the real world, we encourage researchers to think through the ethics of data scraping from publicly available sources to train deep learning models.
+Action recognition is a fundamental computer vision task that involves identifying and classifying human actions within video sequences.
+This technology is critically important in various applications, including security systems, interactive gaming, and healthcare monitoring.
+As social media platforms have grown, TikTok has become a prominent source of diverse human action videos.
+In response, we have developed a specialized foundation model trained on a carefully curated set of 283,582 TikTok video clips.
+This dataset, known as TikTokActions, represents a wide array of human activities and is designed to reflect real-world scenarios and cultural diversities.
+It includes unique, non-standard human actions, filling gaps in existing action recognition datasets and expanding the variety of human activities that current models can recognize.
+Contrasting with the Kinetics dataset[1], which also sources user-generated content but from YouTube, our TikTokActions dataset is composed exclusively of TikTok videos.
+These are characterized by their dynamic and spontaneous nature, often featuring rapid transitions and creative interactions that are less common in YouTube’s typically longer and more structured videos.
+This difference provides unique training challenges, enhancing the ability of models to interpret a broad spectrum of human actions in unpredictable environments.
+We conducted a series of rigorous experiments to evaluate the efficacy of the TikTokActions dataset for pre-training computer vision models on action-related tasks.
+Using the advanced VideoMAE V2 model equipped with ViT-base and ViT-giant backbones[2], we explored the dataset’s capabilities.
+The ViT-base provided a baseline for comparison against other models with similar architectures, while the ViT-giant was used to assess the maximum potential of our dataset.
+Subsequent fine-tuning on well-known datasets such as UCF101[3], HMDB51[4], Kinetics[1], and Something-Something V2[5] revealed that models pre-trained on TikTokActions often performed comparably to, or even better than, those trained solely on these traditional benchmarks or those pre-trained on significantly larger datasets, where enhancements tend to be incremental.
+Additionally, we investigated the relationship between the number of pre-training videos and model performance on downstream tasks.
+By training VideoMAE V2[2] on increasing subsets of our collection, from 1,000 videos up to 6,000, we observed that performance remained robust, suggesting that massive datasets are not always essential for effective pre-training.
+This discovery emphasizes the efficiency of using online, unlabeled videos as a source for pre-training, challenging the prevailing reliance on large-scale video datasets for optimal model performance.
+Prior to deploying such methodologies in practical applications, we urge the research community to consider the ethical implications of using publicly available data for training deep learning models, particularly in terms of data scraping and privacy concerns.
 2 Related Work
 2.1 Human Action Recognition Datasets
-The evolution of action recognition research has been strongly influenced by the availability of benchmark datasets.
-The UCF101[7] dataset, with its diverse collection of human actions, has been a cornerstone for early research in this domain, guiding and benchmarking numerous computer vision innovations applied to action recognition.
-Similarly, the HMDB-51[5] dataset, despite its limited size, has been pivotal in providing a varied set of human actions captured under realistic conditions.
-The Kinetics Human Action Video dataset[4], introduced in 2017, marked a significant milestone in this field.
-With its vast collection of YouTube-sourced videos encompassing 400 distinct human action classes, Kinetics has become an indispensable resource for researchers, setting a higher benchmark and promoting advanced research in action recognition.
-Another noteworthy dataset is the Something-Something[3] collection.
-Its unique focus on human-object interactions captured in real-world scenarios has offered researchers a different perspective, emphasizing the complexity and diversity of everyday actions.
-These datasets, along with several others, have collectively laid the foundation for the current advancements in action recognition, facilitating the development and evaluation of state-of-the-art models and techniques.
-Recognizing the rapid evolution of content on platforms like TikTok and the emergence of diverse, contemporary actions and scenarios, we are motivated to develop a comprehensive dataset highlighting the latest trends and user behaviors specific to new short-video social media platforms.
-Our aim was to provide the community with a dataset with well-defined yet non-standard human actions that complement common actions already represented in current action recognition datasets, thus adding variety to the actions that action recognition models are currently evaluated for.
+The evolution of action recognition research has been significantly shaped by the development and availability of benchmark datasets, which guide and validate advancements in computer vision methodologies.
+The UCF101[3] dataset, known for its diverse array of human actions, has been instrumental in pioneering research within this domain.
+Likewise, the HMDB-51[4] dataset, despite its smaller scale, offers a rich variety of human actions captured under realistic conditions, providing essential challenges for model evaluation.
+Introduced in 2017, the Kinetics Human Action Video dataset [1] became a cornerstone in this field with over 400 distinct human action classes derived from YouTube videos.
+It set a new benchmark, catalyzing further research and development in action recognition technologies.
+Similarly, AVA [6] provides densely labeled video segments that allow for action localization in addition to recognition, offering a deeper understanding of context and action dynamics.
+Emerging datasets such as Ego4D [7], which focuses on first-person video analysis, and HowTo100M [8], which leverages instructional videos for action recognition, reflect the diversifying sources and modalities being explored.
+EPIC-Kitchens [9], ActivityNet [10], and THUMOS [11] offer further variations in complexity and scenarios, pushing the envelope on what models can understand and predict.
+Recent datasets have addressed limitations of earlier collections.
+The Something-Something dataset [5] focuses on human-object interactions within everyday scenarios, offering a unique perspective that highlights the complexity of routine human activities.
+Moments in Time [12] and Multi-Moment Dataset [13] extend this by providing large-scale resources for temporal action understanding, capturing short videos that encompass a wide range of activities.
+The recent Tencent-MVSE dataset [14], another substantial contribution, is designed for multi-modal video similarity evaluation, adding another layer of complexity to action recognition tasks.
+Building upon the diversity and scope of existing datasets, the TikTokActions dataset is meticulously designed to capture the rapidly evolving landscape of social media-driven human interactions.
+Unlike traditional datasets, which often focus on pre-defined, standard action categories, TikTokActions delves into a broader spectrum of spontaneous and culturally varied human behaviors, which are inherently more aligned with contemporary social media trends.
+This alignment is critical as platforms like TikTok foster unique user-generated content that is not only highly dynamic but also reflective of current societal norms and interactions.
+In direct comparison to the Kinetics dataset, which largely aggregates content from YouTube, TikTokActions offers an array of actions driven by the unique constraints and creativity of TikTok’s platform, such as shorter video lengths and a more diverse global user base.
+This distinction is significant as it introduces new challenges in action recognition, particularly in recognizing quick, context-driven actions that are less prevalent in the typically longer-form content of YouTube.
+To substantiate the dataset’s novelty and utility, we conducted a thorough analysis of the action categories present in TikTokActions compared to those in Kinetics and other major benchmarks.
+Our findings reveal that TikTokActions includes a higher proportion of non-standard actions, such as specific dance moves, challenge-based activities, and regional cultural expressions, which are underrepresented in other datasets.
+This inclusion not only enhances the dataset’s coverage of real-world scenarios but also provides a more granular understanding of human actions within the digitally connected world.
+By providing detailed statistics on the distribution of these categories, along with examples in our supplementary materials, we demonstrate the unique value of TikTokActions in filling the existing gaps within the field of action recognition.
+This detailed comparison and analysis affirm that TikTokActions is not just an addition to the plethora of action recognition datasets but a necessary evolution to keep pace with the changing dynamics of human activity in the digital age.
 2.2 Computer Vision Foundation Models
-Foundation models have become essential tools in the field of computer vision.
-The advent of models like Florence[10] showcases the versatility and scope of these models.
-Specifically, Florence encapsulates the progression of foundation models, adeptly offering transitions from broad scene outlines to detailed object-centric views and from static imagery to dynamic video sequences.
+Foundation models in computer vision have revolutionized the way we approach diverse tasks across modalities, offering scalable solutions that extend from image to video understanding.
+The Florence model [15], Specifically, encapsulates the progression of foundation models, adeptly offering transitions from broad scene outlines to detailed object-centric views and from static imagery to dynamic video sequences.
 Additionally, its effectiveness spans various modalities, from RGB imagery to textual data, establishing new standards in tasks like classification, object detection, retrieval, and particularly in action recognition.
-Another significant model in this domain is VideoMAE V2[9].
-Introduced at CVPR 2023, this self-supervised pre-training methodology stands out for its ability to train on datasets exceeding 1 million videos.
+Significant advancements have also been seen with the introduction of VideoMAE V2 [2].
+This self-supervised pre-training methodology stands out for its ability to train on datasets exceeding 1 million videos.
 With its innovative dual masking strategy and progressive training approach, VideoMAE V2 sets new standards in action recognition.
 Empirical evaluations demonstrate its superior performance across multiple benchmarks, emphasizing its effectiveness in real-world scenarios.
-We anticipate that our extensive TikTok dataset will serve as a valuable resource for the continued development and refinement of computer vision foundation models, contributing to the advancement of this rapidly evolving field.
+Further developments include ViViT [16] and TimeSformer [17], which adapt transformer architectures specifically for video, offering improvements in handling the temporal dynamics of action sequences.
+These models demonstrate the potential of transformers in capturing complex, time-related patterns in video data.
+The Quo Vadis model [18] introduced a new approach by integrating the Kinetics dataset to set a performance benchmark, inspiring subsequent models to focus on improving temporal understanding and action recognition capabilities.
+Similarly, Temporal Relational Reasoning in videos [19] provided insights into the importance of temporal dynamics, influencing how subsequent models like Non-local Neural Networks [20] handle long-range dependencies within video sequences.
+A Closer Look at Spatiotemporal Convolutions [21] further explored the convolutional network architectures, focusing on how variations in spatial and temporal modules affect performance on action recognition tasks.
+This study paved the way for more specialized models like the Video Action Transformer Network [22], which introduced an approach to directly model relationships between actions and corresponding contextual features in videos.
+Furthermore, 2D CNNs such as TSN [23], TSM [24], TANet [25], and TDN [26] have been influential by effectively incorporating temporal information, a critical aspect of video understanding.
+The use of 3D CNNs such as I3D [18], R(2+1)D [21], ARTNet [27], and Slow-Fast [28] has further pushed the envelope in spatial-temporal feature integration, enhancing model performance on complex video tasks.
+In the domain of transformers, advancements have continued with models such as MViT [29] and Video Swin Transformer [30] which have introduced ways to integrate spatial and temporal features more effectively, pushing the boundaries of action recognition performance.
+Additionally, UniFormer [31] integrates local and global interactions more efficiently, which is crucial for detailed action understanding in videos.
+The diversity of foundation models extends to multi-modal approaches, where models like Perceiver IO [32] and Data2Vec [33] have been developed to handle various inputs ranging from text to video, showing flexibility and robustness in learning from heterogeneous data sources.
+In the context of enhancing action recognition, SlowFast Networks [28] and X3D [34] introduce novel architectural choices that optimize the processing of temporal variations and the physical intensity of actions.
+These developments underscore the trend towards more specialized and efficient models that are not only capable of high performance but also adaptability across different action recognition scenarios.
+Our TikTokActions dataset, with its focus on spontaneous and diverse user-generated content, provides a rich testing ground for these advanced models.
+By evaluating these models on our dataset, we aim to explore how well they can adapt to and interpret the nuanced, quick-transition actions typical of content found on modern social media platforms.
+The insights gained can help tailor these foundation models to better handle real-world variability and complexity in human actions.
+In our efforts to push the frontiers of action recognition further, we also draw upon the extensive research documented in the Multi-dataset Training of Transformers [35], which explores the robustness gained from training across varied video datasets.
+Such approaches are crucial as they contribute to the generalization capabilities of models, preparing them for deployment in diverse, unstructured environments.
 3 Dataset Construction
@@
 For instance, the use of the hashtag #piano suffices to compile a more extensive collection(see Table  1) of short videos instead of using #pianoplaying.
+Table 1: Example of the expansion in data availability when removing nouns from the hashtag construction process.
 Hashtags Number of Views
@@
 #playingpiano 29.4 Million
-Table 1: Example of the expansion in data availability when removing nouns from the hashtag construction process.
 Figure 1: The hashtag selection and modification process.
@@
 In the domain of human action recognition, benchmarking performance is essential to understand the efficacy of a model.
-This section evaluates the performance of our model, initially pre-trained on the entire collection of 283,582 unlabeled TikTok video clips using VideoMAE V2[9], and subsequently fine-tuned on Kinetics-400 (K400).
+This section evaluates the performance of our model, initially pre-trained on the entire collection of 283,582 unlabeled TikTok video clips using VideoMAE V2[2].
+We employed both the ViT-base and ViT-giant backbones for pre-training to assess the impact of model capacity.
+Subsequently, the model was fine-tuned on Kinetics-400 (K400).
 Following this, the model underwent further fine-tuning on UCF-101[3] and HMDB51[4], adopting a transfer learning approach akin to the experimental setups used in VideoMAE V2 studies.
@@
 VideoMAE V2[2] is a state-of-the-art masked autoencoding framework, notable for its flexibility in backbone selection, ranging from base to billion-level configurations.
-We opted for the ViT-base backbone, which, while less computationally intense than its billion-level counterparts, significantly enhances performance across various video-related tasks, including video action recognition.
+We opted for both the ViT-base and ViT-giant backbones.
+The ViT-base was used to provide a baseline comparison with other models using similarly sized backbones, while the ViT-giant was employed to explore the maximum potential of the dataset, aiming to achieve the highest possible accuracy.
 The generalization capability of VideoMAE V2 pre-trained ViTs[37] as video foundation models has been demonstrated across multiple downstream tasks, highlighting its robustness and adaptability.
 Our training was conducted on 8 NVIDIA A100 40GB GPUs, utilizing the full set of 283,582 unlabeled TikTok videos.
-The pre-training on the ViT-base backbone lasted for 20 days, providing a comprehensive base for developing a robust model capable of understanding a wide array of human actions.
+The pre-training on the ViT-base backbone, which ran for 800 epochs and lasted for 20 days, provided a comprehensive base for developing a robust model capable of understanding a wide array of human actions.
+The ViT-giant backbone was pre-trained under similar conditions, running for 1200 epochs over 30 days, to push the performance boundaries further.
 4.2 Fine-Tuning on Established Datasets
@@
 The first path involved direct fine-tuning on the Kinetics-400 and Something-Something V2 datasets to assess immediate applicability to broad action recognition tasks.
-The results of this direct fine-tuning are detailed in Table 2.
+The results of this direct fine-tuning are detailed in Table 2 and Table 3.
 The second path involved using the model fine-tuned on Kinetics-400 as an intermediary for transfer learning on UCF-101 and HMDB51, leveraging robust pre-training on TikTok videos.
 The outcomes of this transfer learning process are presented in Table 4.
-Dataset Top-1 (%) Top-5 (%)
-Kinetics-400[4] 81.1 94.8
-Something-Something V2[3] 69.2 91.6
-Table 2: Direct Fine-Tuning Results With TikTok Videos Pre-training
-Dataset Top-1 (%) Top-5 (%)
-HMDB-51[5] 83.9 96.7
-UCF-101[7] 97.5 99.6
+Table 2: Results on the Kinetics-400 dataset.
+We scale the pre-training of VideoMAE V2 to billion-level ViT-g model with curated TikTok clips.
+We report the fine-tuning accuracy of multiple view fusion (5×3).
+Both models are pre-trained and fine-tuned at the input of 16×224×224 and sampling stride $\tau$ = 4.
+method pre-train data data size epoch ViT-B ViT-g
+MAE-ST [feichtenhofer2022masked] Kinetics400 0.24M 1600 81.3 -
+VideoMAE V1 [38] Kinetics400 0.24M 1600 81.5 -
+VideoMAE V2 [2] UnlabeledHybrid 1.35M 1200 81.5 87.2
+VideoMAE V2 TikTok 0.28M 800 81.1 -
+VideoMAE V2 TikTok 0.28M 1200 - 85.51
+Table 3: Results on the Something-Something V2 dataset.
+We scale the pre-training of VideoMAE V2 to billion-level ViT-g model with curated TikTok clips.
+We report the fine-tuning accuracy of multiple view fusion (2×3).
+Both models are pre-trained and fine-tuned at the input of 16×224×224 and sampling stride $\tau$ = 2.
+method pre-train data data size epoch ViT-B ViT-g
+MAE-ST [feichtenhofer2022masked] Kinetics400 0.24M 1600 70.8 -
+VideoMAE V1 [38] Sth-Sth V2 0.17M 2400 70.8 -
+VideoMAE V2 [2] UnlabeledHybrid 1.35M 1200 71.2 77.0
+VideoMAE V2 TikTok 0.28M 800 69.2 -
+VideoMAE V2 TikTok 0.28M 1200 - 74.27
 Table 4: Transfer Learning Results From K400 Fine-Tuning Through TikTok Pre-training
-The fine-tuning results highlight the significant potential of using large, unlabeled video collections from platforms like TikTok for advancing human action recognition tasks.
-This approach capitalizes on the rich, varied nature of content available on social media, which, when used for pre-training, enhances the model’s ability to generalize across different datasets and recognition scenarios.
-The robust performance across both direct and transfer learning paths underscores the value of TikTok videos in training models to tackle complex action recognition challenges.
-5 Extrapolating the Required Pre-Training Dataset Size
-In Figure LABEL:fig:accplot, we find a monotone increasing accuracy score versus data volume.
-In the spirit of Floares et al.
-[2], we attempt to extrapolate the number of preprocessed videos for a desired accuracy by curve fitting to the number of data points trained so far.
-Because the accuracy maxes out at 100%, we choose to fit curves that are monotonically increasing with a horizontal asymptote at $11$.
-This makes any CDF of a random variable a viable choice for curve fitting.
-By iterating over the random variables in Scipy [8], we choose the top 10 distributions that fit the accuracy curve.
-Table 4 displays the number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
-We validated these projections by pre-training a model with 9157 data points and fine-tuning the model using the same methods applied in the initial projections.
-The outcome of this validation was notably consistent with our predictive model.
-The model achieved an accuracy of 94.93%, closely aligning with the expected performance as indicated by the inverse CDF projections for this dataset size.
-This validation not only reinforces the reliability of our statistical model but also highlights the practicality of using such predictive techniques in determining dataset size requirements.
-The fact that a dataset size of just 9157 points could yield an accuracy so close to 95% – a benchmark often targeted in action recognition tasks – is a testament to the efficiency of our approach.
-These results suggest that the requirement for massive datasets in the pre-training phase may be reconsidered, especially when computational resources are limited.
-This could mark a significant shift in how datasets are compiled and utilized in machine learning, particularly in fields where data acquisition is a challenge.
-Table 4: Number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
-Model Desired Accuracy
-95% 96% 97% 98% 99%
-t 9157 12388 18323 31892 82574
-Burr XII 9090 12209 17857 30511 76240
-InvGamma 9072 12166 17750 30204 74859
-BetaPrime 9072 12165 17748 30200 74843
-F 9072 12166 17749 30203 74855
-GenPareto 9072 12164 17744 30184 74759
-Lomax 9072 12164 17744 30184 74759
-LogLaplace 9072 12164 17744 30184 74760
-Pareto 9072 12164 17743 30183 74755
-InvWeibull 9043 12101 17596 29772 72915
+Dataset ViT-b Top-1 (%) ViT-g Top-1 (%)
+HMDB-51[4] 83.9 86.08
+UCF-101[3] 97.5 99.05
+Observations on data scaling.
+We studied the influence of pre-training data size on the performance of VideoMAE V2.
+In this experiment, we pre-trained the video models with ViT-B and ViT-g backbones on our TikTok dataset, which contains approximately 0.28M videos.
+The fine-tuning accuracy is detailed in Table 2 for Kinetics-400 and Table 3 for Something-Something V2.
+Despite using approximately five times fewer pre-training samples compared to the UnlabeledHybrid dataset (1.35M videos), our model achieved comparable results.
+Specifically, the ViT-B backbone pre-trained on TikTok data achieved 81.1% on Kinetics-400, which is only 0.4% lower than the 81.5% achieved using the much larger UnlabeledHybrid dataset.
+Similarly, for Something-Something V2, the ViT-B backbone pre-trained on TikTok data achieved 69.2%, compared to 71.2% with the UnlabeledHybrid dataset, showing a difference of 2%.
+These small differences highlight the efficacy of our foundation model on TikTok videos, which did not include any information from the original fine-tuned dataset.
+Observations on model scaling.
+We also explored how different model capacities affect performance.
+The fine-tuning results for models pre-trained with ViT-B and ViT-g backbones are presented in Table 2 and Table 3.
+The ViT-g model, representing a billion-level parameter architecture, consistently demonstrated improved performance over the ViT-B model.
+For example, the ViT-g backbone pre-trained on TikTok data achieved 85.51% on Kinetics-400, compared to 81.1% for the ViT-B backbone, showing an improvement of 4.41%.
+Similarly, for Something-Something V2, compared to the ViT-B backbone, there is an improvement of 5.07%.
+These results suggest that increasing model capacity can further boost performance, validating the potential of large-scale pre-training on diverse video datasets like those from TikTok.
+Observations on transfer learning.
+We further evaluated the performance of our pre-trained models by using Kinetics-400 fine-tuned models for transfer learning on HMDB-51 and UCF-101 datasets.
+The results are summarized in Table 4.
+The ViT-B model achieved 83.9% on HMDB-51 and 97.5% on UCF-101, while the ViT-g model achieved 86.08% on HMDB-51 and 99.05% on UCF-101.
+Overall, our findings highlight the significant potential of using large, unlabeled video collections from platforms like TikTok for advancing human action recognition tasks.
+This approach capitalizes on the rich, varied nature of content available on social media, which, when used for pre-training, enhances the model’s ability to generalize across different datasets and action recognition scenarios.
+The robust performance observed across both direct fine-tuning and transfer learning paths underscores the value of TikTok videos in training models for complex action recognition challenges.
+The TikTok dataset, which is unlabeled and lacks ground truth annotations, still produced promising results, demonstrating the model’s capacity to learn effectively from unstructured video content.
+5 Dataset Size vs. Fine-Tuning Efficacy
+To investigate the impact of pre-training data size on downstream model performance, we varied the size of the pre-training data from 1,000 to 6,000 videos, randomly selected from the TikTokActions dataset.
+Each data size level was used to pre-train the model in three independent runs to measure variance.
+Following pre-training, these models were fine-tuned on the UCF-101[3] dataset.
+The primary metric for evaluating model performance was the accuracy observed on the test set of UCF-101, allowing us to draw conclusions about the relationship between pre-training data size and fine-tuning effectiveness.
+From the result in Figure  3, the volume of pre-trained data from the TikTokActions dataset was incrementally increased, and we observed a positive correlation with the Top-1 accuracy, particularly notable as the volume grew from 1,000 to 3,000 videos.
+However, subsequent increases in data volume yielded a reduced rate of improvement in Top-1 accuracy, indicating diminishing returns despite the larger dataset sizes.
+The Top-5 accuracy continued to show slight improvements, but the incremental benefit also appeared to diminish as the dataset size expanded beyond 3,000 videos.
+These findings suggest an asymptotic behavior in the benefit derived from increasing the volume of pre-training data for our models.
+Figure 3: 100 Epoch Finetune: Final top-1 and top-5 accuracy by amount of pre-training data.
+We demonstrate the impact of the volume of pre-trained TikTok dataset on the fine-tuning accuracy for both Top-1 and Top-5 metrics.
+6 Dataset Size vs. Fine-Tuning Efficacy
+We conducted further experiments to demonstrate the efficacy of the collection in pre-training models for action recognition.
+We used a targeted subset of 6,900 high-quality clips sourced from 73 TikTok hashtags.
+From Figure  4, we can see the pre-trained model achieved a top-1 accuracy of 94.35% and a top-5 accuracy of 99.35% when fine-tuning on the UCF101[3].
+This is remarkably close to the performance of models pre-trained on the UCF101 dataset itself (8155 from 101 classes), which exhibited a top-1 accuracy of 94.50% and a top-5 accuracy of 99.64%.
+The comparable accuracies highlight the mined dataset’s quality and its potential to serve as a substantial pre-training resource, even with fewer and more focused training samples.
+The model trained from scratch, without the benefit of such a rich dataset, performed significantly poorer, reinforcing the value of pre-training on high-quality, domain-specific data.
+This outcome attests to the robustness and generalization capabilities of models pre-trained on the curated dataset.
+Figure 4: We compare the top-1 and top-5 accuracy of models pre-trained on TikTok actions, pre-trained from scratch, and pre-trained on UCF101 when fine-tuned to UCF101.
 7 Discussion
@@
 8 Ethics
-This research adheres to strict ethical guidelines regarding the use of online video data, specifically videos collected from TikTok.
-All videos were legally accessed via the TikTok Research API, and only those videos where the user granted download permission were included.
-Our data collection process is in full compliance with the TikTok Research API Terms of Service.
-As an additional layer of security, we delete the data after training.
-We place a high priority on privacy and copyright issues.
+This study has been approved under the University of Hawaii Institutional Review Board (IRB) under protocol number 2024-00396.
+Furthermore, this research adheres to strict ethical guidelines regarding the use of online video data, specifically videos collected from TikTok.
+All videos were legally accessed via the TikTok Research API, and our data collection process is in full compliance with the TikTok Research API Terms of Service.
+Importantly, only those videos where the user granted explicit download permission were included in our training process.
+Equally as importantly, we deleted all of the data after model training as an additional layer of security.
+We place a high priority on user privacy.
 To protect individual privacy, we have implemented face obfuscation techniques in the video clips included in our paper.
@@
 Instead, we focus on training a foundation model, which we only use to identify action patterns, not personal biometric features.
-Ethical Considerations in Online Video Research: The use of online video data poses unique challenges and ethical considerations, particularly regarding privacy, informed consent, and the potential risks associated with the use of such data.
+Genderal Ethical Considerations for Online Video Research: The use of online video data poses unique challenges and ethical considerations, particularly regarding privacy, informed consent, and the potential risks associated with the use of such data.
 Our study emphasizes the importance of adhering to ethical principles that safeguard participant interests.
-We have made significant efforts to minimize potential harm by anonymizing the data and not sharing the scraped dataset.
-Although techniques like pixelating faces can be circumvented and are not foolproof, they represent our commitment to reducing identifiable information.
-Furthermore, we advocate for the inclusion of clearer guidelines in the terms of service on video platforms, which would allow users to explicitly exclude their videos from research use.
+We have made significant efforts to minimize potential harm by anonymizing the data and not sharing the curated dataset.
+We recommend that researchers adhere to such privacy protections.
+Furthermore, we advocate for the inclusion of clearer guidelines in the terms of service on video platforms, which would allow users of the platforms to explicitly exclude their videos from research use.
 This approach does not replace the need for informed consent but is a step towards enhancing ethical practices in online video research.
 We aim to foster a transparent and informed discussion about the use of online video data in computational research.
+We also argue that the use of the models plays a critical role in deciding the ethical standards that are required.
+Predicting human actions such as running, swimming, or playing piano from videos is a common task that is unlikely to lead to harm, especially given how common action recognition computer vision research is.
+By contrast, if the AI were used to predict sensitive properties about the end user such as mental health diagnoses, income, political party, or anything else along those lines, then the standard for explicit consent and ethical practices would rise to a much higher level [39].
+[…]
 Acknowledgements
@@
 The AWS Cloud Credit for Research program has also played a crucial role in providing computational resources for our work.
-We are grateful for the funding provided by the National Institutes of Health under the NIH Director’s New Innovator Award (award DP2-EB035858).
 We used ChatGPT to edit the grammar of our manuscript and to re-phrase sentences that were originally worded unclearly.
 However, all contents in this manuscript are original ideas and analyses conducted by the authors.
-Appendix A Appendix
+Appendix A Sample Clips
 In the appendix, we will present additional insights into our dataset, specifically focusing on visual and statistical aspects.
@@
 These elements combined will give a more complete picture of the dataset’s richness.
-Appendix B Sample Clips
 In Figure 5, we present the extensive diversity of the TikTokActions dataset, which captures a wide range of human activities tagged across multiple domains such as fitness, sports, and dance.
@@
 The presence of such varied content provides a rich learning ground for self-supervised learning models to extract and learn a comprehensive set of features from human actions, gestures, and interactions, aiding in the enhancement of action recognition capabilities.
-Figure 3: Example classes from the TikTokActions dataset.
-Appendix C Dataset Composition
+Figure 5: Example clips from the TikTokActions dataset.
+Appendix B Dataset Composition
 Our dataset offers a diverse array of video counts across different hashtags.
@@
 Figure 7: Video counts in different ranges, highlighting the dataset’s variety and the volume of data available for model training.
-Appendix D Dataset Duration
+Appendix C Dataset Duration
 In the composition of the TikTokActions dataset, a key focus was on ensuring that each video clip is sufficiently long to capture a complete action, while also maintaining a manageable length for analysis.
@@
 The following table details the distribution of video durations within our dataset, post these adjustments:
+Table 5: Distribution of video durations in the TikTokActions dataset. The data reflects the adherence to the minimum and maximum duration thresholds for clip selection.
 Duration Range (seconds) Number of Videos
@@
 10+ 172123
-Table 5: Distribution of video durations in the TikTokActions dataset. The data reflects the adherence to the minimum and maximum duration thresholds for clip selection.